Gellie is twice as old as Harry. If 4 years is added to Gellie's age and 2 is subtracted from Harry's age, Gellie will be three times as old as Harry. Find their present ages.
Answer by ankor@dixie-net.com(22740)

Gellie is twice as old as Harry.
g = 2h
:
If 4 years is added to Gellie's age and 2 is subtracted from Harry's age, Gellie will be three times as old as Harry.
g + 4 = 3(h-2)
g + 4 = 3h - 6
g = 3h - 6 - 4
g = 3h - 10
replace g with 2h, find h, you should be able to do this now



Find their present ages
: